REGULARITY OF HOLOMORPHIC CORRESPONDENCES AND APPLICATIONS TO THE MAPPING PROBLEMS

XIAOJUN HUANG

Abstract. We study the regularity results of holomorphic correspondences. As an application, we combine it with certain recently developed methods to obtain the extension theorem for proper holomorphic mappings between domains with real analytic boundaries in the complex 2-space.